# LETS-C: Leveraging Language Embedding for Time Series Classification

### v1

@@ Abstract
Recent advancements in language modeling have shown promising results when applied to time series data. In particular, fine-tuning pre-trained large language models (LLMs) for time series classification tasks has achieved state-of-the-art (SOTA) performance on standard benchmarks. However, these LLM-based models have a significant drawback due to the large model size, with the number of trainable parameters in the millions. In this paper, we propose an alternative approach to leveraging the success of language modeling in the time series domain. Instead of fine-tuning LLMs, we utilize a language embedding model to embed time series and then pair the embeddings with a simple classification head composed of convolutional neural networks (CNN) and multilayer perceptron (MLP). We conducted extensive experiments on well-established time series classification benchmark datasets. We demonstrated LETS-C not only outperforms the current SOTA in classification accuracy but also offers a lightweight solution, using only 14.5% of the trainable parameters on average compared to the SOTA model. Our findings suggest that leveraging language encoders to embed time series data, combined with a simple yet effective classification head, offers a promising direction for achieving high-performance time series classification while maintaining a lightweight model architecture.
@@ 1 Introduction
Time series classification Bagnall et al. (2017); Abanda et al. (2019); Ismail Fawaz et al. (2019) has gained significant attention in recent years due to its wide-ranging applications in various domains, such as finance Passalis et al. (2017), healthcare Lipton et al. (2016), and activity recognition Yang et al. (2015). The increasing availability of time series data has driven the need for efficient and accurate classification methods. Recent advancements in natural language processing (NLP) and large language models (LLMs) have shown strong promises in language modeling Achiam et al. (2023), particularly in capturing temporal dependencies within sequential data. Inspired by this success, researchers have explored extending these techniques to the time series domain by fine-tuning pre-trained LLMs Zhou et al. (2024); Jin et al. (2023), achieving state-of-the-art (SOTA) performance on well-established benchmarks for tasks including classification and forecasting.
@@
However, the use of LLMs for time series classification comes with a significant drawback due to their large model size. These models often have billions of parameters, making them computationally expensive and limiting their usage in resource-constrained environments Bommasani et al. (2021). At training time, fine-tuning partially frozen pre-trained LLMs also often involves millions of trainable parameters Zhou et al. (2024). To address this challenge, we propose an alternative approach to leverage the success of language modeling in the time series domain. In particular, we propose a novel approach called LETS-C (Language Embeddings for Time Series Classification) that leverages language embeddings, also known as text embeddings, for time series classification. To the best of our knowledge, this work is the first to explore the potential of language embeddings in time series analysis, specifically classification, and demonstrate SOTA performance.
@@
LETS-C combines language embeddings with a simple yet effective classification head composed of convolutional neural networks (CNNs) and a multilayer perceptron (MLP). By projecting time series data using language embedding models, we capture the intricate patterns and dependencies present in the temporal data. The embeddings and time series are then fed into the classification head, which learns to discriminate between different classes. Through extensive experiments on a standard benchmark containing 10 datasets across various domains Bagnall et al. (2018), we demonstrated that LETS-C outperforms 20 baselines including the previous SOTA method. Moreover, LETS-C is significantly more efficient, using much less trainable parameters than the previous SOTA method.
Our main contributions are as follows:
@@
  • Language Embeddings for Time Series: We introduce LETS-C, the first work to leverage language embeddings (or text embeddings) for time series analysis, specifically for classification tasks.
@@
  • State-of-the-Art Performance: LETS-C achieves SOTA performance in classification accuracy on a well-established benchmark containing 10 datasets across different domains, surpassing 20 baselines.
@@
  • Lightweight: LETS-C is significantly more efficient, achieving higher accuracy with much fewer trainable parameters (14.5%) compared to the existing SOTA method.
In addition, we conducted comprehensive analysis to along different aspects to showcase the effectiveness of LETS-C:
  • Text Embeddings Models: LETS-C with different text embedding models consistently outperforms previous SOTA with much fewer trainable parameters, further validating the effectiveness of our approach.
@@
  • Time Series Embeddings: We revealed the advantage of text embeddings on time series, showing the embeddings of time series from the same class are more similar than the ones from different classes, explaining the boost in classification accuracy
  • Trade-off between Accuracy and Model Size: LETS-C retains a high percentage of accuracy even when the model size shrinks considerably, making it even more computationally efficient without compromising too much on model accuracy.
The remainder of this paper is organized as follows: Section 2 briefly summarizes related work. Section 3 details the structure of the proposed approach. In Section 4, we describe the experimental setup, benchmark results, and additional findings from comprehensive evaluations. Finally, Section 5 discusses our findings and Section 6 outlines limitations and future research directions. Due to space limitations, a more extensive discussion of the experimental setup and results is provided in the Appendix.
@@ 2 Related Work
In this section, we review the related works in three key areas: time series classification, the application of language models to time series data, and text embeddings.
@@
Time Series Classification Time series classification has been an active research area for decades. We include an extended review in Appendix A due to space limit. Early methods focused on distance-based approaches Abanda et al. (2019), such as Dynamic Time Warping (DTW) Berndt & Clifford (1994) and distance kernels with Support Vector Machines (SVMs) Kampouraki et al. (2008). Others extracted features and used linear or tree-based classifiers like eXtreme Gradient Boosting (XGBoost) Chen & Guestrin (2016). Later, deep learning-based approaches, including Convolutional Neural Networks (CNNs) Wu et al. (2022a); Zhao et al. (2017), Multilayer Perceptron (MLP) Zhang et al. (2022), and Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) Lai et al. (2018), gained popularity. These models can learn complex patterns and handle long sequences. Recently, Transformer-based models Vaswani et al. (2017) have been adapted from Natural Language Processing (NLP) to the time series domain Zhou et al. (2022); Wu et al. (2021); Zhou et al. (2021). They use self-attention mechanisms to model long-range dependencies. However, these more complex models often have larger sizes and higher computational costs, especially for training.
@@
Language Models for Time Series The success of language modeling in NLP and LLMs has inspired researchers to harness LLMs in the time series domain. Comprehensive surveys Zhang et al. (2024); Jiang et al. (2024) have offered valuable insights into the integration of LLMs in time series analysis, highlighting key methodologies, challenges, and future directions.  Gruver et al. (2023); Xue & Salim (2023) enabled pre-trained LLMs to generate time series forecasts through prompting. Further,  Yu et al. (2023) delved into the potential of LLMs for generating explainable forecasts of financial time series.  Jin et al. (2023) introduced Time-LLM, which focuses on learning to project time series into the language embedding space and directly using pre-trained LLM for time series forecasting tasks. More importantly, recent work OneFitsAll by Zhou et al. (2024) has shown promising results by fine-tuning models like GPT Radford et al. (2019) on time series tasks. OneFitsAll achieved SOTA performance compared across various time series tasks including classification. Inspired by these successes, we propose to explore another form of language modeling, text embeddings, in the time series domain. We discover that this direction leads to new SOTA performance on time series classification tasks, with the benefit of much fewer trainable parameters than OneFitsAll.
@@
Text Embeddings Text embeddings have played a crucial role in NLP. These embeddings map words or sentences into a dense vector space, capturing semantic and syntactic information. Various text embedding techniques have been proposed, ranging from word-level embeddings like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) to contextualized embeddings obtained from pre-trained language models such as BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019b). In the time series domain, some works have proposed unsupervised methods for learning time series embeddings Franceschi et al. (2019); Tonekaboni et al. (2021). However, the availability of large-scale datasets in the time series domain is generally more limited compared to those in the NLP domain, making learning time series embedding from scratch more challenging compared to text embeddings. To our best knowledge, we are the first to leverage the well-trained text embeddings from the NLP domain for time series classification.
@@ 3 Methodology
Given a time series classification dataset $\mathcal{D}=\{(\mathbf{x_{i}},y_{i})_{i=1}^{N}\}$, where $\mathbf{x_{i}}$ is a multivariate time series sample, and $y_{i}\in\{1,2,\dots,C\}$ is the corresponding class label, the goal is to learn a classifier that accurately predicts the class label $\hat{y}_{i}$ for each time series. As illustrated in Figure 1, we propose LETS-C framework that harnesses text embeddings for time series classification tasks. Specifically, we 1) initially preprocess the time series data to standardize it, then 2) subsequently generate text embeddings from the standardized time series, 3) fuse embeddings with the time series data, and finally 4) feed the fused representation to a classification head that consists of CNNs and MLP. The choice of a simple classification head is intentional, we aim to test the hypothesis that the text embeddings of the time series provide sufficiently powerful representations for effective classification.
@@
Figure 1: Left: Conventional text embedding. Right: Our proposed LETS-C framework normalizes time series and formats them to tokenize each digit separately. It then embeds the time series into the embedding space, fuses the embeddings and time series together via element-wise addition, and employs a simple classification head consisting of a CNN and an MLP for classification. In this framework, the only elements trained are the lightweight CNN model and the MLP head.
@@
Preprocessing To ensure consistent scales across all model inputs, each feature dimension of time series $\mathbf{x_{i}}$ is min-max normalized to the range $[0,1]$ based on the minimum and maximum feature values of each dimension across the training data.
@@
Text Embedding of Time Series It is crucial to carefully format the preprocessed time series into strings before using text embeddings, as the tokenization of numerical strings can significantly affect the embeddings.  Liu & Low (2023) has shown that tokenization impacts a model’s arithmetic abilities, with commonly used subword tokenization methods like Byte Pair Encoding (BPE) arbitrarily subdividing numbers, causing similar numbers to appear very differently. To mitigate this, we adopted a digit-space tokenization strategy, as suggested by Gruver et al. (2024), where each digit is spaced, commas are added to separate time steps, and decimal points are omitted for fixed precision. For instance, a series with a precision of two decimal places, such as 0.645, 6.45, 64.5, 645.0, would be formatted as "6 4 , 6 4 5 , 6 4 5 0 , 6 4 5 0 0" prior to tokenization. This method ensures separate tokenization of each digit, preserving numerical integrity and enhancing pattern recognition in language models.
@@
Next, we utilized OpenAI’s text-embedding-3-large model OpenAI (2024) to embed the formatted time series into the embedding space. This model was selected for several reasons: it is highly ranked on the Massive Text Embedding Benchmark (MTEB) leaderboard MTEB (2024); Muennighoff et al. (2022), known for its effectiveness in a variety of downstream tasks such as text search and sentence similarity; it supports a high maximum token length of 8191, accommodating our time series datasets; and it offers a high-dimensional vector space of 3072 dimensions. This large dimensionality captures a broad spectrum of temporal features, yet the model allows for truncation to reduce dimensions as needed for specific applications without substantial loss of semantic information Kusupati et al. (2022). This capability to truncate dimensions is particularly advantageous for optimizing computational efficiency while maintaining robust performance in downstream applications, aligning well with our goal of a lightweight framework.
In particular, we take each dimension of $\mathbf{x_{i}}$ and generate the corresponding text embedding. Thus, we transform $\mathbf{x_{i}}\in\mathbb{R}^{d\times l_{x}}$ into embeddings $\mathbf{e_{i}}\in\mathbb{R}^{d\times l_{e}}$, where $d$ is the multivariate dimension, $l_{x}$ is the length of the time series, and $l_{e}$ is the length of the embedding. Alternatively, each time series can be divided into separate patches where each patch has a corresponding text embedding. However, this alternative is not scalable and did not yield significant performance benefits empirically. Note that the embedding computation in LETS-C is a one-time pass, contrasting with the persistent computational cost by models such as OneFitsAll which rely on partially frozen transformer-based components.
To validate that text embeddings are suitable for time series classification, we compared the cosine similarities of text embeddings of time series from the same and different classes. Our results show consistently higher intra-class similarity than inter-class. This highlights the effectiveness of text embeddings in time series tasks (see Section 4.3).
Fusing Embedding and Time Series Next, we perform an element-wise addition of the embedding to the preprocessed time series, while preserving the maximum length between the two. Specifically, $\mathbf{x_{i}}\in\mathbb{R}^{d\times l_{x}}$ are added with embeddings $\mathbf{e_{i}}\in\mathbb{R}^{d\times l_{e}}$. If $l_{x}$ and $l_{e}$ do not match, we pad the shorter one with zeros to align their sizes before addition, resulting in a combined representation in $\mathbb{R}^{d\times max(l_{x},l_{e})}$. The direct addition of embeddings to the time series data is analogous to the use of positional embeddings in NLP, which enriches the sequence representation with additional structural and contextual information. As a result, it enables both time series and text embeddings to contribute to the learning process, thereby enhancing the model’s ability to effectively interpret complex temporal patterns (Section 4.3 show alternative fusion approaches are less effective).
@@
Lightweight Classification Head Lastly, we pair the fused embedding and time series with a simple classification head composed of 1D CNNs and an MLP for time series classification. The output from CNNs are flattened and fed through the final MLP head, which outputs a vector of the probabilities of each time series class. Hyperparameter search determines the number of convolutional blocks in the CNN, the number of linear layers in the MLP, and the utilization of batch normalization, dropout, activation functions, and pooling operations. With a simple classification head, our model is lightweight and requires much less trainable parameters compared to the existing SOTA method built on transformers, as benchmarked in Section 4.2.
4 Experiments
We provide a brief introduction to the datasets, baselines, evaluation metrics, and model implementation details in Section 4.1. In Section 4.2, we compare our approach against various baselines—including classical methods and MLP-, CNN-, RNN-, and Transformer-based techniques, as well as recent SOTA LLM-based methods—highlighting LETS-C’s performance and computational costs. Further insights into ablation studies of LETS-C, generalization of our method across different embedding types, the power of the embeddings and its similarities across different time series are explored through additional analyses in Section 4.3.
4.1 Experimental Setup

### v2

# LETS-C: Leveraging Text Embedding for Time Series Classification

@@ Abstract
Recent advancements in language modeling have shown promising results when applied to time series data. In particular, fine-tuning pre-trained large language models (LLMs) for time series classification tasks has achieved state-of-the-art (SOTA) performance on standard benchmarks. However, these LLM-based models have a significant drawback due to the large model size, with the number of trainable parameters in the millions. In this paper, we propose an alternative approach to leveraging the success of language modeling in the time series domain. Instead of fine-tuning LLMs, we utilize a text embedding model to embed time series and then pair the embeddings with a simple classification head composed of convolutional neural networks (CNN) and multilayer perceptron (MLP). We conducted extensive experiments on a well-established time series classification benchmark. We demonstrated LETS-C not only outperforms the current SOTA in classification accuracy but also offers a lightweight solution, using only 14.5% of the trainable parameters on average compared to the SOTA model. Our findings suggest that leveraging text embedding models to encode time series data, combined with a simple yet effective classification head, offers a promising direction for achieving high-performance time series classification while maintaining a lightweight model architecture.
LETS-C: Leveraging Text Embedding for Time Series Classification
Rachneet Kaur  Zhen Zeng  Tucker Balch  Manuela Veloso J.P. Morgan AI Research {rachneet.kaur, zhen.zeng, tucker.balch, manuela.veloso}@jpmorgan.com
@@ 1 Introduction
Time series classification Bagnall et al. (2017); Abanda et al. (2019); Ismail Fawaz et al. (2019) has gained attention due to its wide-ranging applications in various domains, such as finance Passalis et al. (2017), healthcare Lipton et al. (2016), and activity recognition Yang et al. (2015). The increasing availability of time series data has driven the need for efficient and accurate classification methods. Advances in natural language processing (NLP) and large language models (LLMs) Achiam et al. (2023) have demonstrated strong promises in modeling sequential data Achiam et al. (2023). Inspired by this, researchers have explored applying LLMs to time series via prompting Gruver et al. (2024); Liu et al. (2024); Merrill et al. (2024); Chu et al. (2024) or fine-tuning pre-trained LLMs Jin et al. (2023); Zhou et al. (2024), achieving state-of-the-art (SOTA) performance on well-established benchmarks for tasks including classification and forecasting.
@@
However, LLMs for time series classification face a major drawback—their large size, often with billions of parameters, making them computationally expensive and impractical in resource-limited settings Bommasani et al. (2021). Fine-tuning partially frozen pre-trained LLMs still requires millions of trainable parameters Zhou et al. (2024). To mitigate this, we propose an alternative approach to leverage the success of language modeling in the time series domain. In particular, we propose a novel approach, LETS-C (Leveraging Text Embeddings for Time Series Classification), which utilizes off-the-shelf text embedding models instead of fine-tuning LLMs for time series classification. To the best of our knowledge, this work is the first to explore the potential of text embeddings in time series analysis, specifically classification, and demonstrate SOTA performance.
@@
LETS-C combines text embeddings with a simple yet effective classification head composed of convolutional neural networks (CNNs) and a multilayer perceptron (MLP). By projecting time series data using text embedding models, we capture the intricate patterns and dependencies present in the temporal data. The embeddings and time series are then fed into the classification head, which learns to discriminate between different classes. Through extensive experiments on a well-established benchmark containing time series datasets from various domains, we demonstrated that LETS-C outperforms 27 baselines including the previous SOTA method. Moreover, LETS-C is significantly more efficient, using much fewer trainable parameters than the previous SOTA. Our key contributions are:
@@
  • Text Embeddings for Time Series: We introduce LETS-C, the first work to leverage text embeddings for time series analysis, specifically for classification tasks.
@@
  • State-of-the-Art Performance: LETS-C achieves SOTA performance in classification accuracy on a well-established benchmark containing time series datasets across different domains, surpassing 27 baseline models.
@@
  • Computationally Lightweight: LETS-C is significantly more efficient, achieving higher accuracy while using much fewer trainable parameters (14.5%) than the existing SOTA method.
In addition, we conducted comprehensive analyses to showcase the effectiveness of LETS-C:
@@
  • Built-in Discriminative Power of Time Series Embeddings: We demonstrate the advantage of text embeddings for time series, showing that embeddings of time series from the same class are more similar than those from different classes, explaining the boost in classification accuracy.
  • Generalization Across Various Text Embedding Models: LETS-C with different text embedding models consistently outperforms previous SOTA with much fewer trainable parameters, further validating the effectiveness of our approach.
  • Optimizing Model Size with Minimal Accuracy Loss: LETS-C retains a high percentage of accuracy even as the classification model size shrinks considerably, enhancing computational efficiency with minimal impact on performance.
@@ 2 Related Work
We review the related works in three key areas: time series classification, the application of language models to time series data, and text embeddings. See Appendix A for an extended review.
Time Series Classification
@@
Time series classification has been an active research area for decades. Early methods focused on distance-based approaches Abanda et al. (2019), such as Dynamic Time Warping Berndt and Clifford (1994) and distance kernels with Support Vector Machines  Kampouraki et al. (2008). Others used feature extraction with classifiers like eXtreme Gradient Boosting (XGBoost) Chen and Guestrin (2016) and LightGBM Ke et al. (2017). Deep learning-based approaches, including CNNs Wu et al. (2022a); Zhao et al. (2017), MLPs Zhang et al. (2022a), and Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) Lai et al. (2018), later gained popularity for learning complex patterns and handling long sequences. More recently, Transformer-based models Vaswani et al. (2017) have been adapted from NLP to time series Nie et al. (2023); Zhou et al. (2022); Zhang et al. (2022b); Eldele et al. (2021); Wu et al. (2021); Zhou et al. (2021); Koval et al. (2024), leveraging self-attention for long-range dependencies. Additionally, unsupervised representation learning methods pre-train models with masked time series modeling to minimize reconstruction error, followed by fine-tuning for downstream tasks like classification Franceschi et al. (2019); Tonekaboni et al. (2021); Yue et al. (2022); Eldele et al. (2021); Zerveas et al. (2021); Goswami et al. (2024). However, these complex models often require larger sizes and higher computational costs, particularly for training.
Language Models for Time Series
@@
The success of LLMs in NLP has inspired their application in time series analysis. Surveys Zhang et al. (2024); Jiang et al. (2024) provide insights into key methodologies, challenges, and future directions. Recent studies explore integrating time series and language, including structured time series-text modeling Khadanga et al. (2019); Deznabi et al. (2021), natural language descriptions of time series Murakami et al. (2017); Jhamtani and Berg-Kirkpatrick (2021); Fons et al. (2024), and various applications Li et al. (2024a); Drinkall et al. (2024); Kawarada et al. (2024). Pre-trained LLMs have been used for time series forecasting via prompting Gruver et al. (2024); Liu et al. (2024); Cao et al. (2023), while Yu et al. (2023) explored their potential for generating explainable financial forecasts. Time-LLM Jin et al. (2023) maps time series to the language embedding space, enabling LLM-based forecasting Yang et al. (2021). More importantly, recent work, OneFitsAll Zhou et al. (2024) achieved SOTA performance on various time series tasks by fine-tuning LLMs like GPT Radford et al. (2019). In contrast, we propose leveraging text embeddings instead of directly using LLMs for time series analysis. Our approach establishes new SOTA performance on time series classification tasks with significantly fewer trainable parameters, making it a more efficient alternative.
Text Embeddings
@@
Text embeddings are crucial in NLP, mapping words or sentences into dense vector spaces to capture semantic and syntactic information. Text embedding techniques range from word-level embeddings like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) to contextualized embeddings from pre-trained models like BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019b). In time series, unsupervised methods have been proposed for learning embeddings Franceschi et al. (2019); Eldele et al. (2021); Zerveas et al. (2021); Yue et al. (2022); Sun et al. (2023); Goswami et al. (2024). However, large-scale datasets are scarcer in time series than in NLP, making it more challenging to learn embeddings from scratch. To our knowledge, we are the first to leverage well-trained text embeddings from NLP for time series classification.
@@ 3 Methodology
Given a time series classification dataset $\mathcal{D}=\{(\mathbf{x_{i}},y_{i})_{i=1}^{N}\}$, where $\mathbf{x_{i}}$ is a multivariate time series sample, and $y_{i}\in\{1,2,\dots,C\}$ is the corresponding class label, the goal is to learn a classifier that accurately predicts the class label $\hat{y}_{i}$ for each time series. As illustrated in Figure 1, we propose LETS-C framework that harnesses text embeddings for time series classification tasks. Specifically, we 1) initially preprocess the time series data to normalize it, then 2) subsequently generate text embeddings from the normalized time series, 3) fuse embeddings with the time series data, and finally 4) feed the fused representation to a classification head that consists of CNNs and MLP. The choice of a simple classification head is intentional, we aim to test the hypothesis that the text embeddings of the time series provide sufficiently powerful representations for effective classification.
@@
Figure 1: Left: Conventional text embedding. Right: Our proposed LETS-C framework normalizes time series and formats them to tokenize each digit separately. It then embeds the time series, fuses the embeddings with the original series via element-wise addition, and uses a simple CNN-MLP classification head to perform classification. Only the lightweight CNN and MLP head are trained. For illustration, we show a single-dimensional time series $\mathbf{x_{i}}\in\mathbb{R}^{l_{x}}$, transformed into an embedding vector $\mathbf{e_{i}}\in\mathbb{R}^{l_{e}}$.
Preprocessing
@@
To ensure consistent scales across all model inputs, each feature dimension of time series $\mathbf{x_{i}}$ is min-max normalized to the range $[0,1]$ based on the minimum and maximum feature values of each dimension across the training data.
Text Embedding of Time Series
@@
It is crucial to carefully format the preprocessed time series into strings before using text embeddings, as the tokenization of numerical strings can significantly affect the embeddings.  Liu and Low (2023) has shown that tokenization impacts a model’s arithmetic abilities, with commonly used subword tokenization methods like Byte Pair Encoding (BPE) arbitrarily subdividing numbers, causing similar numbers to appear very differently. To mitigate this, we adopted a digit-space tokenization strategy, as suggested by Gruver et al. (2024), where each digit is spaced, commas are added to separate time steps, and decimal points are omitted for fixed precision. For instance, a series to be formatted with a precision of two decimal places, such as 0.645, 6.45, 64.5, 645.0, would be converted to "6 4 , 6 4 5 , 6 4 5 0 , 6 4 5 0 0" prior to tokenization. This method ensures separate tokenization of each digit, preserving numerical integrity and enhancing pattern recognition in language models.
@@
Next, we utilized the text-embedding-3-large model OpenAI (2024) to embed the formatted time series into the embedding space. It is important to note that we are using only the text embedding model, unlike the LLMs used in previous works Jin et al. (2023); Zhou et al. (2024); Gruver et al. (2024). This model was selected for several reasons: it is highly ranked on the Massive Text Embedding Benchmark (MTEB) leaderboard Muennighoff et al. (2022); MTEB (2024), known for its effectiveness in a variety of downstream tasks such as text search and sentence similarity; it supports a high maximum token length of 8191, accommodating our time series datasets; and it offers a high-dimensional vector space of 3072 dimensions. This large dimensionality captures a broad spectrum of temporal features, additionally the model allows for truncation to reduce dimensions as needed for specific applications without substantial loss of semantic information Kusupati et al. (2022). This capability to truncate dimensions is particularly advantageous for optimizing efficiency while maintaining robust performance, aligning well with our goal of a lightweight framework.
In particular, we generate a text embedding for each dimension of $\mathbf{x_{i}}$, transforming $\mathbf{x_{i}}\in\mathbb{R}^{d\times l_{x}}$ into embeddings $\mathbf{e_{i}}\in\mathbb{R}^{d\times l_{e}}$. Here, $d$ is the number of dimensions in the multivariate time series, $l_{x}$ is the series length, and $l_{e}$ is the embedding length. Specifically, each dimension of $\mathbf{x_{i}}$ (of length $l_{x}$) is first converted into a string, resulting in $d$ separate strings. Each string is then independently embedded using an embedding model, as studies Nie et al. (2023); Zhou et al. (2024); Goswami et al. (2024) suggest that channel-wise modeling is an effective strategy for multivariate time series. Our approach differs slightly depending on the type of embedding model used: 1) Proprietary model: For OpenAI’s text-embedding-3-large, the API directly returns an embedding vector of size $l_{e}$ for each string; 2) Open-weight models: We extract the last token’s embedding from the model’s final hidden states to represent the entire sequence, effectively transforming the dimension from $l_{x}$ to $l_{e}$. This approach captures the final contextual representation by condensing the sequence into a single embedding derived from the last token. Embeddings for each dimension are then combined into a $d\times l_{e}$ matrix, effectively transforming the input from $d\times l_{x}$ to $d\times l_{e}$. This transformation preserves independent contextual representations for each dimension while ensuring a consistent embedding structure. Note that the embedding computation in LETS-C is a one-time pass, then the computed embeddings of the time series can be stored and reused for further training, in contrast to the persistent computational cost caused by fine-tuning parts of LLMs. To validate the suitability of off-the-shelf text embeddings for time series classification, we compared cosine similarities between text embeddings of time series from the same and different classes. Intra-class similarity was consistently higher, demonstrating their effectiveness (see Section 5.2).
Fusing Embedding and Time Series
Next, we fuse the embedding with the preprocessed time series using element-wise addition, applying zero padding for dimensional consistency. The time series embedding vector serves as a feature representation of temporal patterns present in the entire time series, and adding the raw time series integrates both information sources. This approach is well-established, particularly in ResNet He et al. (2016), where raw data is combined with feature representations via shortcut connections. Additionally, fusing embeddings from different modalities is widely supported Manzoor et al. (2023); Guo et al. (2019); Poria et al. (2018); Baltrušaitis et al. (2018); Jabri et al. (2016), with element-wise addition being a common and effective method for combining multimodal embeddings. This fusion allows both time series and text embeddings to contribute meaningfully to the final representation, improving the model’s overall performance. As shown in Section 5.3, using either the text embedding or the raw time series alone, as well as alternative fusion methods to addition, is less effective. While element-wise addition is an empirical design choice in our architecture, our experiments demonstrate that it achieves SOTA performance while maintaining minimal model complexity compared to alternatives such as concatenation, or fusion.
Lightweight Classification Head
@@
Lastly, we pair the fused time series representation with a simple classification head composed of 1D CNNs and an MLP for time series classification. The output from CNNs are flattened and fed through the final MLP head with a softmax activation, which outputs a vector of the probabilities of each time series class. As detailed in the Appendix I, hyperparameter search determines the number of convolutional blocks in CNN, the number of linear layers in MLP, and the use of batch normalization, dropout, activation, and pooling. With a simple classification head, our model is lightweight and requires much fewer trainable parameters compared to the existing SOTA built on transformers, as detailed in Section 5.1.
4 Experimental Protocol and Details
Datasets and Evaluation Metrics